. Solve the following system of equations.




3x + 4y = 4
2x + y = 6



2. Solve the following system of equations.




2x - 3y = 13
5x + 2y = 4


3. The Kraft Co. manufactures computer chips at a variable cost of $4 per chip and sells them for $10 each. If the fixed cost is $12,000 per month, what is the number of chips they would need to produce at the break-even point?


4. The Sunshine Bakery sells pies at a fixed price of p dollars per pie. The total number of pies demanded daily, D, is related to the price, p, in dollars by the equation:




D = -10p + 200


On the other hand, the daily supply of pies, S, is related to the price, p, per pie by the equation:




S = 15p - 50


Determine the equilibrium price of pies; that is, the price at which the supply, S, and demand, D, are equal.


Answer by ewatrrr(24785)


Hi
3x + 4y = 4
2x + y = 6
3x + 4y = 4
-8x -4y = -24
-5x = -20
x = 4 and y = -2 (y = 6-2*4)
2x - 3y = 13
5x + 2y = 4
2x - 3y = 13
15/2x + 3y = 12/2 |mulitplying 2nd EQ thru by 3/2
9.5x = 19
x = 2 and y = -3 (y = -5/2x + 2) |10-6 = 4
3) $6 profit per bag $12000/$6 = 2000bags

4) -10p + 200 = 15p - 50
250 = 25p
$10 = p